# Theoretically motivated dark electromagnetism as the origin of relativistic MOND

### 14 Dec 2023 v1

@@ I Introduction
As we explain in Section VI of the paper, our proposal for MOND as dark electromagnetism also explains Verlinde’s entropic criterion Verlinde for MOND. Keeping this connection with Verlinde in mind, in Section III we review Verlinde’s proposal for motivating MOND from entropy considerations.
Since MOND has a distance dependence in the acceleration as $1/R$, the associated MOND potential is logarithmic. This is in principle consistent with the source being a surface square-root mass density $\sim\sqrt{M}/R^{2}$, as if the MOND dynamics were taking place effectively in two spatial dimensions. This is consistent with the logarithmic form for the Green’s function of the Laplace equation in 2D, as we recall in Section IV. Note however that this surface density of square-root mass does not have a well-defined limit as $R\rightarrow 0$ (it diverges as $R^{-1/2}$), whereas the volume density of square root mass defined using the effective distance does have a well-defined limit which is finite. Also, we would not like to modify the structure of the left-hand side of the Poisson equation, and this is consistent with proposing MOND as the non-relativistic limit of the $U(1)_{DEM}$ gauge theory.
@@
II A brief review of MOND and relativistic MOND
@@
We make mention though of an ongoing related research of great importance: testing the law of gravitation in GAIA DR3 wide binaries Gaia ; g2 ; g3 . A large number of such binaries are known in the solar neighbourhood of the Milky Way, and these have orbital radii ranging from about 200 AU to 30000 AU. The orbital acceleration crosses the critical MOND value $a_{M}$ for radii around 1000 AU, transiting from the Newtonian regime (relatively low radii) to the alleged MOND regime (large radii). Around 2000 AU onwards, the measured acceleration should disagree with Newtonian prediction, if MOND is right. The analyses by Chae Chae1 ; Chae2 and by Hernandez Hernandez1 ; Hernandez2 shows that Newtonian gravitation is obeyed in the not so wide binaries, but breaks down for larger separations. Banik et al. disagree Banik . See however, Chae’s critical response to Banik, and the responses of Lasenby, Boyle, and especially of Hernandez, after the recent OSMU23 lecture by Banik OSMUBanik . See also the recent rebuttal by Hernandez and Chae HC . To our understanding, the conclusion of Chae, and of Hernandez, that Newtonian gravitation breaks down below the critical acceleration $a_{M}$, is correct. It is remarkable that wide binaries have the same critical acceleration scale $a_{M}$ as spiral galaxies do: there is no a priori reason for this to be so, unless the fifth force does indeed exist and begins to dominate gravitation below $a_{M}$. This anomaly in wide binaries cannot be explained by dark matter; therefore wide binaries are the likely smoking gun which will discriminate MOND from dark matter.
III A brief review of Verlinde’s entropic derivation of MOND
III.1 Introduction
@@
III.2 The flattening of galaxy rotation curves
@@
III.3 Entropy as a criterion for a phase transition
III.3.1 Entropy increase when a bit traverses a horizon
@@
III.3.2 Entropy of empty dS space
@@
III.3.3 Entropy reduction of dS space due to the addition of matter
@@
III.3.4 The missing mass problem in entropic terms
@@
III.3.5 Newtonian gravity in terms of surface densities
@@
III.4 The elastic phase of emergent gravity
III.4.1 Elastic moduli in terms of gravitational parameters
@@
III.4.2 A derivation of the Tully–Fisher relation
@@
III.4.3 Apparent dark matter in terms of baryonic matter
@@
IV MOND as gravity in 2+1 dimensions
IV.1 Green’s functions for the Laplacian in two dimensions
IV.2 The plane $\mathbb{R}^{2}$
Given the Euclidean metric on the plane
${\rm d}s^{2}={\rm d}r^{2}+r^{2}{\rm d}\varphi^{2},$ (37)
The solution reads
$G(r)=\left[A-\Theta(r)\right]\ln r+B$ (41)
$G(r)=-\Theta(r)\ln r$ (42)
which, for $r>0$, simplifies to
$G(r)=-\ln r$ (43)
IV.3 The sphere $\mathbb{S}^{2}$
Given the standard round metric on the unit sphere $\mathbb{S}^{2}$,
${\rm d}s^{2}={\rm d}\theta^{2}+\sin^{2}\theta\,{\rm d}\varphi^{2},$ (44)
The change of variables $x=\cos\theta$ reduces (47) to
$(1-x^{2})g^{\prime\prime}(x)-2xg^{\prime}(x)=\delta(x)$ (48)
where we have set $g(x)=G(\theta)$. Eq. (48) is solved by
where $A,B$ arbitrary integration constants. We may set $A=B=0$ to obtain
$G(\theta)=-\ln\theta+\ldots$ (51)
IV.4 Hyperbolic space $\mathbb{H}^{2}$
The metric on hyperbolic space $\mathbb{H}^{2}$ is
${\rm d}s^{2}={\rm d}\varrho^{2}+\sinh^{2}\varrho\,{\rm d}\varphi^{2}$ (52)
We apply the change of variables $x=\cosh\varrho$ to obtain
$(x^{2}-1)g^{\prime\prime}(x)+2xg^{\prime}(x)=-\delta(x),$ (56)
$G(\varrho)\simeq-\ln\varrho+\ldots$ (58)
@@ V Dark electromagnetism as the origin of relativistic MOND
This is the contribution to apparent dark matter from the Coulomb part of the potential energy. If we assume that each of the three vector components also contribute in equal measure, we get that the total apparent dark matter is $5.164\times\sqrt{5/3}=$ 5.16 times ordinary matter. This agrees well with the standard LCDM model according to which dark matter to ordinary matter ratio is about $5.3$.
@@
$S=\frac{c^{3}}{16\pi G}\int d^{4}x\sqrt{-g}\;[R-2\Lambda]+\int d^{4}x\sqrt{-g}\;{\cal L}_{matter}-\frac{1}{16c}\int d^{4}x\sqrt{-g}\;F_{ij}F^{ij}+A\int d^{4}x\sqrt{-g}D_{i}J^{i}$ (72)
@@
Milgrom m2 writes that: “… one may conjecture that the MOND-cosmology connection is such that local gravitational physics would take exactly the deep-MOND form in an exact de Sitter universe. This is based on the equality of the symmetry groups of $dS^{4}$ and of the MOND limit of the Bekenstein-Milgrom formulation BM both groups being $SO(4,1)$. The fact that today we see locally a departure from the exact MOND-limit physics–i.e., that the interpolating functions have the form they have, and that $a_{0}$ is finite and serves as a transition acceleration–stems from the departure of our actual space-time from exact $dS^{4}$ geometry: The broken symmetry of our space-time is thus echoed in the broken symmetry of local physics.” Our proposal that $U(1)_{DEM}$ is the remnant broken symmetry after breaking of $SU(2)_{R}\times U(1)_{YDEM}$ is entirely in support of this conjecture of Milgrom.
@@ VIII.1 Critical acceleration
These results could have important implications in early universe cosmology. In particular, it could be that the electroweak symmetry breaks when the acceleration of a quasi-deSitter expanding universe falls below a critical value (assuming the inflation-like phase ends at the electroweak scale).
@@ VIII.3 Advantages in considering dark electromagnetism
7. We are able to derive Verlinde’s results for apparent dark matter and entropy criterion for MOND
@@
The dark photon - the gauge boson that mediates quantised DEM, could be thought of as dark matter. Its detection in the laboratory may however be beyond current technology. Same could be said about dark electromagnetic waves, though they could well be the early dark radiation Lasenby ; Karwal that has been proposed as one possible solution to the Hubble tension. Even though the dark photon can be called the sought for dark matter, what is noteworthy is that the associated DEM field is MONDian in character, and we have a newly predicted fifth force mimicking dark matter, but not a new fermionic elementary particle as dark matter. From the point of view of fundamental physics, this difference (i.e. is dark matter fermionic or bosonic) is significant. It decides whether there are only four fundamental forces, or more than four.

### 14 Mar 2024 v3

@@ I Introduction
The cold dark matter paradigm is a cornerstone of modern cosmology. Dark matter plays a central role in explaining structure formation in the early universe, as well as in explaining the observed anisotropies of the cosmic microwave background. Dark matter is also successful in explaining gravitational lensing of large scale structures, and the observed baryon acoustic oscillations. Anomalous velocity dispersions in galactic clusters are also accounted for by proposing the existence of dark matter, and with some challenges, galaxy rotation curves can also be explained by assuming a specific density profile for the dark matter. Thus a single assumption, that of a prevailing collisionless and gravitationally attractive fluid with mass density about 5-6 times that of ordinary matter, can account for a wide range of cosmological phenomena. A new elementary particle, which is beyond the standard model of particle physics, is considered the most likely origin of this dark fluid.
And yet, the greatest challenge for dark matter is that no such particle has been experimentally detected in the laboratory so far. And this in spite of a very large number of proposed theoretical candidates, and an intense experimental search over some four decades and through nearly fifty different experiments. Of course this does not imply that such a particle will never be detected, but it does make one ask if there could be an alternative explanation of cosmic phenomena in which dark matter is not required, but say the law of gravitation [Newton’s law as well as general relativity] might be getting modified on large scales. It is not impossible that such a modification to gravitation perfectly mimics the proposed dark fluid. Thus what we call dark matter could in fact be modified gravity in disguise. It is this idea which is explored in the present paper, starting from a first principles proposal for modified gravity, which we call dark electromagnetism.
The most successful proposal for a modified law of gravitation, as an alternative to dark matter, is Milgrom’s Modified Newtonian Dynamics [MOND]. This is a phenomenological proposal in which, for accelerations far smaller than a certain low critical value, the law of gravitation (in the non-relativistic limit) changes from inverse square to a $1/R$ law. MOND does very well at explaining galaxy rotation curves; whereas on cluster scales and cosmic scales it has modest success and also faces serious challenges. MOND’s major challenge is that it lacks a first principles theoretical explanation: what are the convincing theoretical reasons (independent of observations) which might compel us to modify general relativity, and that too in such a way that MOND is an inevitable consequence of the modified theory in its non-relativistic limit?
Our recent proposal for a unified theory of interactions provides strong evidence for a theoretical origin of MOND, in the form of a fifth force. Surprisingly, this unified theory does not provide any dark matter particle candidate, thus favouring modified gravity over dark matter. This prediction also makes the theory eminently falsifiable - it will be rule out by a laboratory detection of dark matter. The present paper describes the origin of this fifth force (dark electromagnetism) and how it serves as the origin of MOND. The theory is briefly reviewed in Section II, which is followed by a brief review of MOND in Section III.
As we explain in Section VI of the paper, our proposal for MOND as dark electromagnetism also explains Verlinde’s entropic criterion Verlinde for MOND. Keeping this connection with Verlinde in mind, in Section IV we review Verlinde’s proposal for motivating MOND from entropy considerations.
Section V is central to the paper, and explains how dark electromagnetism mimics MOND on various cosmic scales. Related empirical aspects of dark electromagnetism are discussed in Section VII. Section VIII critiques our findings and we also comment on the current challenges faced by MOND.
II Theoretical origin of dark electromagnetism
@@
Since MOND has a distance dependence in the acceleration as $1/R$, the associated MOND potential is logarithmic. This is in principle consistent with the source being a surface square-root mass density $\sim\sqrt{M}/R^{2}$, as if the MOND dynamics were taking place effectively in two spatial dimensions. This is consistent with the logarithmic form for the Green’s function of the Laplace equation in 2D, as we review in the Appendix in Section VIII. Note however that this surface density of square-root mass does not have a well-defined limit as $R\rightarrow 0$ (it diverges as $R^{-1/2}$), whereas the volume density of square root mass defined using the effective distance does have a well-defined limit which is finite. Also, we would not like to modify the structure of the left-hand side of the Poisson equation, and this is consistent with proposing MOND as the non-relativistic limit of the $U(1)_{DEM}$ gauge theory.
@@
III A brief review of MOND and relativistic MOND
@@
We make mention though of an ongoing related research of great importance: testing the law of gravitation in GAIA DR3 wide binaries Gaia ; g2 ; g3 . A large number of such binaries are known in the solar neighbourhood of the Milky Way, and these have orbital radii ranging from about 200 AU to 30000 AU. The orbital acceleration crosses the critical MOND value $a_{M}$ for radii around 1000 AU, transiting from the Newtonian regime (relatively low radii) to the alleged MOND regime (large radii). Around 2000 AU onwards, the measured acceleration should disagree with Newtonian prediction, if MOND is right. The analyses by Chae Chae1 ; Chae2 and by Hernandez Hernandez1 ; Hernandez2 shows that Newtonian gravitation is obeyed in the not so wide binaries, but breaks down for larger separations. Banik et al. disagree Banik . See however, Chae’s critical response to Banik Chae2 , and the responses of Lasenby, Boyle, and especially of Hernandez, after the recent OSMU23 lecture by Banik OSMUBanik . See also the recent rebuttal by Hernandez and Chae HC . To our understanding, the conclusion of Chae, and of Hernandez, that Newtonian gravitation breaks down below the critical acceleration $a_{M}$, is correct. It is remarkable that wide binaries have the same critical acceleration scale $a_{M}$ as spiral galaxies do: there is no a priori reason for this to be so, unless the fifth force does indeed exist and begins to dominate gravitation below $a_{M}$. This anomaly in wide binaries cannot be explained by dark matter; therefore wide binaries are the likely smoking gun which will discriminate MOND from dark matter.
We also note that not all researchers are agreed on the presence of a universal critical acceleration scale in galaxies. In their analysis of 193 disk galaxies from the SPARC and THINGS databases, Rodrigues et al. conclude the absence of a fundamental acceleration scale in galaxies Rodrigues . Whereas, in an accompanying paper Gaugh et al. perform a Bayesian analysis on galaxy rotation curves from the SPARC database and find strong evidence for a characteristic acceleration scale Gaugh . See also the critical analysis in Chan and references therein. Our outlook is that these analyses are a part of an ongoing debate, and for the purpose of our present theoretical discussion we will assume that a fundamental acceleration scale does exist on galactic scales.
Dark matter is a cornerstone assumption in contemporary cosmology, supported by evidence from galaxy rotation curves, gravitational lensing of bullet cluster, CMB fluctuations and baryon acoustic oscillations (BAO), and the formation of large-scale structures. Therefore it is important to assess how MOND fares vis a vis these aspects, as an alternative to dark matter. MOND does spectacularly well with galaxy rotation curves, being able to predict the rotation curve once the baryonic mass distribution of a galaxy is known from observations. In this regard it does better than the cold dark matter hypothesis, where a rotation curve has to be first known from observations and then a CDM distribution is assumed so as to fit the curve of velocities. MOND also provides a natural explanation for the Tully-Fisher relation, which is a challenge for CDM. MOND does not adequately account for cluster velocity dispersions, where CDM certainly does better. It has however been suggested that MOND’s shortfall on cluster scales could be accounted for by the missing baryons which are currently unaccounted for in clusters.
CDM has a definite upper hand when it comes to formation of large scale structures, CMB anisotropies, BAO and gravitational lensing. Here one is in need of a convincing relativistic MOND which generalises general relativity at low accelerations and can then be convincingly applied in cosmology. The present proposal for dark electromagnetism is a step in that direction. For a detailed recent review of the status of MOND in astrophysics and cosmology, we refer the reader to Banik and Zhao BZ .
IV A brief review of Verlinde’s entropic derivation of MOND
IV.1 Introduction
@@
IV.2 The flattening of galaxy rotation curves
@@
IV.3 Entropy as a criterion for a phase transition
IV.3.1 Entropy increase when a bit traverses a horizon
@@
IV.3.2 Entropy of empty dS space
@@
IV.3.3 Entropy reduction of dS space due to the addition of matter
@@
IV.3.4 The missing mass problem in entropic terms
@@
IV.3.5 Newtonian gravity in terms of surface densities
@@
IV.4 The elastic phase of emergent gravity
IV.4.1 Elastic moduli in terms of gravitational parameters
@@
IV.4.2 A derivation of the Tully–Fisher relation
@@
IV.4.3 Apparent dark matter in terms of baryonic matter
@@ V Dark electromagnetism as the origin of relativistic MOND
The introduction of a characteristic acceleration ($a_{M}$) related to the Hubble constant implies variability over time, suggesting that galaxies at different redshifts would exhibit distinct rotation curves. This aspect has been discussed carefully for instance in the MOND review by Milgrom MilgromS (see the subsection on ‘The significance of the MOND acceleration constant’ therein, as well as the reference to Milgrom2017 where high redshift rotation curves are discussed in the context of MOND). Milgrom concludes that there are strong observational constraints on the variation of $a_{M}$ with cosmic time, and a value of $4a_{M}$ at $z\sim 2$ is essentially ruled out, hence excluding a dependence such as $a\propto(1+z)^{3/2}$.
@@
This is the contribution to apparent dark matter from the Coulomb part of the potential energy. If we assume that each of the three vector components also contribute in equal measure, we get that the total apparent dark matter is $5.164\times\sqrt{5/3}=$ 5.16 times ordinary matter. This agrees well with the standard LCDM model according to which dark matter to ordinary matter ratio is about $5.3$. The assumption that the vector components contribute in equal measure as the Coulomb part is a reasonable one because these considerations are being applied on the scale of the entire universe, including at high redshifts. Therefore relativistic motions must be taken into account, and as a result the ‘magnetic part’ of the four-potential is expected to be as significant as the Coulomb part.
@@
$S$ $=\frac{c^{3}}{16\pi G}\int d^{4}x\sqrt{-g}\;[R-2\Lambda]+\int d^{4}x\sqrt{-g}\;{\cal L}_{matter}$
$-\frac{1}{16c}\int d^{4}x\sqrt{-g}\;F_{ij}F^{ij}+A\int d^{4}x\sqrt{-g}D_{i}J^{i}$ (50)
@@
The solution to this equation determines the DEM field, which then enters the Einstein equations as a source, as an alternative to dark matter. This source term is the potential energy of the DEM field, and it represents the enhanced gravitational interaction amongst baryons, instead of having to invoke dark matter to provide the sought for additional gravitational effects. The right hand side of Eqn. (45) above is an illustration of this claim.
@@
Milgrom m2 writes that: “… one may conjecture that the MOND-cosmology connection is such that local gravitational physics would take exactly the deep-MOND form in an exact de Sitter universe. This is based on the equality of the symmetry groups of $dS^{4}$ and of the MOND limit of the Bekenstein-Milgrom formulation BM both groups being $SO(4,1)$. The fact that today we see locally a departure from the exact MOND-limit physics–i.e., that the interpolating functions have the form they have, and that $a_{0}$ is finite and serves as a transition acceleration–stems from the departure of our actual space-time from exact $dS^{4}$ geometry: The broken symmetry of our space-time is thus echoed in the broken symmetry of local physics.” Our proposal that $U(1)_{DEM}$ is the remnant unbroken symmetry after breaking of $SU(2)_{R}\times U(1)_{YDEM}$ is entirely in support of this conjecture of Milgrom.
A few further remarks about the proposed action principle in Eqn. (50) are in order. This form of the action is assumed to come into play after the electroweak symmetry breaking around a TeV scale. That is the epoch at which classical spacetime, general relativity and dark electromagnetism (i.e. relativistic MOND) emerge. This emergence is expected to give the same physical results as standard big bang cosmology, with cold dark matter exchanged for dark electromagnetism. Prior to this emergence, cosmology is governed by the unified $E_{8}\times E_{8}$ theory, above the TeV scale. Since cosmological data are not yet available at such high energy scales, there is no contradiction between the octonionic theory and cosmology of the very early universe.
@@ VIII.1 Critical acceleration
These results could have important implications in early universe cosmology. In particular, it could be that the electroweak symmetry breaks when the acceleration of a quasi-deSitter expanding universe falls below a critical value (assuming the inflation-like phase ends at the electroweak scale). See however the critical analysis in this regard, by Unruh and Weiss Unruh and by Hill Hill .
@@ VIII.3 Advantages in considering dark electromagnetism
7. We are able to derive Verlinde’s results for apparent dark matter and entropy criterion for MOND.
@@
The dark photon - the massless gauge boson that mediates quantised DEM, could be thought of as dark matter. Its detection in the laboratory may however be beyond current technology. Same could be said about dark electromagnetic waves, though they could well be the early dark radiation Lasenby ; Karwal that has been proposed as one possible solution to the Hubble tension. Even though the dark photon can be called the sought for dark matter, what is noteworthy is that the associated DEM field is MONDian in character, and we have a newly predicted fifth force mimicking dark matter, but not a new fermionic elementary particle as dark matter. From the point of view of fundamental physics, this difference (i.e. is dark matter fermionic or bosonic) is significant. It decides whether there are only four fundamental forces, or more than four.
After this paper was published, we received a private communication from Davi C. Rodrigues, co-author of Ref. [41]. They have expressed concern about our remark on their work, on p. 13 above (towards the end of Section III). We quote their comment:
“We clarify here that Ref. [41] did a Bayesian analysis, and computed the $5\sigma$ uncertainty interval of $a_{0}$ for each galaxy; concluding that there is a high level of tension between these results (more than 5 or 10 $\sigma$). Ref. [42] did a quite different analysis, based on the reduced $\chi^{2}$. Criticisms raised by [42] were promptly answered in the reply Rodrigues:2018lvw . The debate about the relevance of the priors in this context, likewise the development of analysis improvements, has continued after these works.”
We thank Davi Rodrigues for this clarification.
@@
IX APPENDIX: MOND as gravity in 2+1 dimensions
IX.1 Green’s functions for the Laplacian in two dimensions
IX.2 The plane $\mathbb{R}^{2}$
Given the Euclidean metric on the plane
${\rm d}s^{2}={\rm d}r^{2}+r^{2}{\rm d}\varphi^{2},$ (58)
The solution reads
$G(r)=\left[A-\Theta(r)\right]\ln r+B$ (62)
$G(r)=-\Theta(r)\ln r$ (63)
which, for $r>0$, simplifies to
$G(r)=-\ln r$ (64)
IX.3 The sphere $\mathbb{S}^{2}$
Given the standard round metric on the unit sphere $\mathbb{S}^{2}$,
${\rm d}s^{2}={\rm d}\theta^{2}+\sin^{2}\theta\,{\rm d}\varphi^{2},$ (65)
The change of variables $x=\cos\theta$ reduces (68) to
$(1-x^{2})g^{\prime\prime}(x)-2xg^{\prime}(x)=\delta(x)$ (69)
where we have set $g(x)=G(\theta)$. Eq. (69) is solved by
where $A,B$ arbitrary integration constants. We may set $A=B=0$ to obtain
$G(\theta)=-\ln\theta+\ldots$ (72)
IX.4 Hyperbolic space $\mathbb{H}^{2}$
The metric on hyperbolic space $\mathbb{H}^{2}$ is
${\rm d}s^{2}={\rm d}\varrho^{2}+\sinh^{2}\varrho\,{\rm d}\varphi^{2}$ (73)
We apply the change of variables $x=\cosh\varrho$ to obtain
$(x^{2}-1)g^{\prime\prime}(x)+2xg^{\prime}(x)=-\delta(x),$ (77)
$G(\varrho)\simeq-\ln\varrho+\ldots$ (79)